3. (20 points) Streiber Publishing Company, an all-equity firm, generates perpetual earnings before interest and taxes of $2.5 million per year. Its after-tax, all-equity discount rate is 20%. The companys tax rate is 34%. a. What is the value of Streiber Publishing? b. If it adjusts its capital structure to include $600,000 of debt, what is the value of the firm? Explain any difference in your answers. What assumptions are you making when you are valuing Streiber? c. d.
Answer #1

EBIT = $ 2.5 million per year

Ke = 20%

Tax rate, T = 34%

A. Value of Firm = EBIT(1-T)/Ke = 2.5(1-0.34)/0.2

Value of the firm = $ 8.25 million

B. Now the company introduces debt of $ 600,000 in its capital structure.

The value of Firm can be obtained using the MM approach

large V_L = V_U + TB

+ TB K.

2.5(1 0.34) 0.2 +0.34 *0.6

8.25 + 0.204

Value of levered firm = $ 8.454 million

C. The difference in the market value is due to addition of debt in the capital structure of the firm. The cost of debt is relatively cheaper than the cost of equity and the firm get a tax shield due to the debt. As a resultant the rate of return increases which improves the value of Firm. This is termed as Financial Leverage.

Difference in MV = $ 0.204 million

D. All the assumptions of MM approach

1. Perfect capital market

2. Homogenous risk class

3. Homogenous expectation.

Please contact if having any query thank you.